OpenStudy (anonymous):

Factor: 20a^2-13a-21=0 Please show work. Thanks :)

OpenStudy (anonymous):

This is a bit tricky because you have so many possible factor pairs. You're looking for something like (Ax+B)(Cx+D) It would multiply out to: \[ACx^2+(BC+AD)x+BD\] So in your case AC=20, BD=-21, (BC+AD)=-13

OpenStudy (anonymous):

Now look at the factor pairs that can give you -21: 1,21; 3,7 where one of the numbers is negative. Factor pairs for 20 are: 1,20; 2,10; 4,5 So you can look for combinations of BC+AD that add to -13.

OpenStudy (anonymous):

One trick is to look at factors of 420 (20x210) that are 13 apart: 1,2,3,4,5,6,7,10,12,15,20,21,28,30,35,42,60,70,84,105,140,210,420 I probably should have listed them as factor pairs but 15,28 works

OpenStudy (anonymous):

We know that AC=20 and AD=-28 so A=4,C=5,D=-7,B=3 (4a+3)(5x-7)

OpenStudy (anonymous):

The "trick statement should have read ..."(20x21)"